OpenStudy (anonymous):

can somebody help? Evaluate each polynomial for the given value of the variable. 2x2-4x-1; x= 5

OpenStudy (anonymous):

@narissa

OpenStudy (bloomlocke367):

just plug 5 into for x and solve.. do you know how to do that?

OpenStudy (bloomlocke367):

@Danny.Mcintire7

OpenStudy (anonymous):

no I dont

OpenStudy (anonymous):

2(5)2-4(5)-1=

OpenStudy (anonymous):

@bloomlocke367

OpenStudy (bloomlocke367):

yea.. \[2\times (5)^2-4\times (5)-1=?\] do you know how to do that?

OpenStudy (anonymous):

50-20-1=29

OpenStudy (bloomlocke367):

there you go! XD

OpenStudy (bloomlocke367):

\(\Huge\mathbb{\color{lime}{CORRECT!!!!!!!}}\)

OpenStudy (anonymous):

would the answer just be 29 or 50-20-1=29

OpenStudy (bloomlocke367):

the answer is just 29 XD

OpenStudy (anonymous):

K thanks!:]

OpenStudy (bloomlocke367):

\(\scr\Huge\color{purple}{You're~Welcome!}\)

OpenStudy (bloomlocke367):

\(\Huge\color{pink}{\heartsuit}\color{#2efec8}{\heartsuit}\color{pink}{\heartsuit}\color{#2efec8}{\heartsuit}\color{pink}{\heartsuit}\color{#2efec8}{\heartsuit}\color{pink}{\heartsuit}\color{#2efec8}{\heartsuit}\color{pink}{\heartsuit}\color{#2efec8}{\heartsuit}\color{pink}{\heartsuit}\) \(\bbox[ 5pt, #f5a9f2, border: dashed #f781d8 5pt]{\scr\Huge\color{#01dfa5}{BloomLocke367}}\) \(\Huge\color{pink}{\heartsuit}\color{#2efec8}{\heartsuit}\color{pink}{\heartsuit}\color{#2efec8}{\heartsuit}\color{pink}{\heartsuit}\color{#2efec8}{\heartsuit}\color{pink}{\heartsuit}\color{#2efec8}{\heartsuit}\color{pink}{\heartsuit}\color{#2efec8}{\heartsuit}\color{pink}{\heartsuit}\)